# Why Reinforcement Fine-Tuning Enables MLLMs Preserve Prior Knowledge Better: A Data Perspective

### v3

@@ Abstract
Post-training algorithms such as Supervised Fine-Tuning (SFT) and Reinforcement Fine-Tuning (RFT) are widely used to adapt multimodal large language models to downstream tasks. While effective at task adaptation, their impact on prior knowledge remains unclear. In this paper, we introduce jigsaw puzzles as a novel task absent from existing pretraining corpora and systematically study the behavior of SFT and RFT on open-source multimodal model, Qwen2.5-VL series. Our experiments reveal a sharp trade-off: SFT enables rapid task acquisition but leads to catastrophic forgetting, whereas RFT learns more slowly but maintains prior knowledge. We study this phenomenon through learning dynamics by examining both the magnitude and direction of how training data influence prior knowledge. Our analysis shows that RFT mainly reinforces correct samples naturally aligned with the base model’s probability landscape, leading to weaker interference with prior knowledge. Moreover, training on RFT-simulated rollouts, which exert a small magnitude of influence and are well aligned in direction to prior knowledge, allows SFT to preserve prior knowledge better while rapidly learning new tasks. These findings suggest that distribution of training data, rather than algorithmic differences, plays a central role in forgetting, and highlight RFT’s potential for stable continual learning in multimodal large language models.
@@ 1 Introduction
In the era of large models, two primary post-training methods, i.e., Supervised Fine-Tuning (SFT) (wei2021finetuned) and Reinforcement Fine-Tuning (RFT) (DeepSeek-R1; ouyang2022training), have emerged for enhancing model performance on domain-specific tasks. These methods have been pivotal in enabling multimodal large language models (MLLMs) to learn specific downstream tasks, follow human instructions, and acquire reasoning capabilities, yielding impressive results. However, current post-training practices primarily focus on performance improvement for specific downstream tasks, while overlooking the potential impact of fine-tuning algorithms on the model’s pre-existing knowledge. This oversight raises concerns about the model’s ability to retain and apply prior knowledge.
@@
To this end, this paper investigates how post-training algorithms, specifically SFT and RFT, affect the retention of prior knowledge when large models are trained to learn entirely novel knowledge or tasks that were absent during pretraining. In particular, we will focus our study on MLLMs, which understand both vision and language. To establish a challenging and genuinely novel task for testing, we introduce jigsaw puzzles as the target task for learning, as in Fig. 1. Through preliminary experiments, we observe that existing state-of-the-art MLLMs, including GPT-4o (hurst2024gpt), fail to solve even simple 3x3 jigsaw puzzles, indicating that this task represents a novel problem not covered by current pretraining corpora. Thus, the jigsaw puzzles can serve as a fair and meaningful task for evaluating the impact of post-training algorithms, i.e., SFT and RFT, on prior knowledge.
@@
We conduct systematic fine-tuning experiments using both standard SFT and RFT, i.e., GRPO (shao2024deepseekmath), on open-sourced Qwen2.5-VL (Qwen2.5-VL) series. Interestingly, we find that SFT can master novel tasks with solely hundreds or thousands of training steps, while RFT requires several tens of thousands of training steps to successfully solve jigsaw puzzles and achieves similar accuracy as SFT. This finding suggests that large-scale RFT can teach model to solve tasks that base model is completely unable to handle. In a sense, this shows that RFT can push the model beyond its original capability boundary. In addition to performance on novel tasks, we observe that SFT incurs severe forgetting of previous knowledge, with substantial performance drops across diverse benchmarks, especially on tasks with similar output formats as jigsaw puzzles. In contrast, RFT, by leveraging reward-driven credit assignments for the simulated rollouts, can master the novel jigsaw puzzles while maintaining decent performance on prior tasks.
@@
The distinct phenomenon observed between SFT and RFT naturally raises a question: Why SFT incurs catastrophic forgetting while RFT does not? While both algorithms increase the likelihood of correct responses, RFT adaptively reweights the likelihood of rollout with reward, whereas SFT uniformly increases the likelihood of static human annotations. Inspired by this, we collect the correct rollouts during RFT and use them as the supervised data for SFT. Surprisingly, we find that SFT trained on these correct rollouts not only acquires novel knowledge quickly but also preserves prior knowledge better. This suggests that the construction of fine-tuning data, rather than the training algorithm itself, is a key factor in knowledge forgetting.
@@
Furthermore, we provide a new perspective based on learning dynamics (ren2024learning), which links the likelihood change of prior knowledge to the gradient induced by an individual training example , on understanding this distinct forgetting behavior by analyzing the magnitude and direction of how training data influence prior knowledge. We first observe that SFT data without reasoning trajectories usually interfere more with prior knowledge, as verified with a much larger norm of empirical neural tangent kernel (eNTK) between SFT data and prior knowledge. While datasets with reasoning trajectories, such as reasoning trajectories generated by GPT-4o and collected during RFT, usually exhibit a smaller norm of eNTK and less forgetting of prior knowledge, implying that introducing explicit reasoning can help alleviate knowledge forgetting.
@@
As for the direction of interference, we find that data with reasoning trajectories generated by GPT-4o typically belong to high-perplexity regions of the base model. In contrast, data collected during RFT are naturally generated from regions where the base model already assigns a moderate probability. This suggests that pretraining has already shaped certain linguistic regions by accident that are well-aligned with novel tasks, while remaining compatible with prior knowledge. Importantly, according to learning dynamics, the influence of training on one example over the likelihood of another example is symmetric: increasing the likelihood of has the same marginal effect on as increasing has on . Therefore, when RFT discovers and reinforces such hidden linguistic regions shaped during pre-training, it degrades less the likelihood of prior knowledge . Crucially, such regions are difficult to identify during the stage of dataset construction for SFT, but are accessible through RFT’s active exploration within linguistic space. This highlights RFT as an effective algorithm for stable novel knowledge acquisition in MLLMs without suffering from catastrophic forgetting. Formally, our contributions are three-fold:
@@
We show that large-scale RFT can solve unseen tasks while preserving prior competencies. Moreover, SFT trained on RFT-generated rollouts can match RFT’s performance while markedly reducing catastrophic forgetting, underscoring the central role of data construction.
@@ 2 Related Works
Jigsaw Puzzles. Jigsaw puzzles has long been a popular self-supervised task in the computer vision community, aimed at learning visual representations (Unsupervised-Learning-of-Visual-Representations-by-Solving-Jigsaw; Domain-Generalization-by-Solving-Jigsaw-Puzzles) by spatial reasoning and part-whole understanding. Recently, this task has been repurposed for probing weak spot of MLLMs: Jigsaw-Puzzles shows that leading MLLMs perform far behind than human performance. The contemporary work Jigsaw-R1 (Jigsaw-R1) solves jigsaw puzzles with RFT, achieving much better performance. Collectively, these works mainly treat jigsaw puzzles as pretext task for representation learning or test benchmark for MLLMs. However, we employ jigsaw puzzles to investigate how post-training algorithms affect the forgetting behavior of MLLMs.
@@
Reinforcement Fine-Tuning in MLLMs. Inspired by the success of RFT in large language models (DeepSeek-R1; ouyang2022training), recent work has applied RFT to MLLMs. Among them, MM-Eureka finds that RFT can achieve better out-of-distribution generalization performance than SFT. Meanwhile, RFT is also employed for perception-centric tasks (liu2025visual; VLM-R1; Seg-Zero), still conferring notable gains in generalization and robustness. Concurrently, Jigsaw-R1 (Jigsaw-R1) introduced RFT to the novel task of jigsaw puzzles but achieved limited accuracy. Building on this direction, we extend RFT training to tens of thousands of steps to enable deeper exploration, yielding substantial gains on jigsaw puzzles.
@@
Catastrophic Forgetting. Early work (mccloskey1989catastrophic; ratcliff1990connectionist) showed that even minimal sequential training on disjoint data can cause rapid “catastrophic forgetting” (CF). Existing strategies to mitigate CF fall into three categories: (i) Regularization-based methods (Overcoming-catastrophic-forgetting-in-neural-networks; Continual-Learning-Through-Synaptic-Intelligence; li2017learning) constrain updates to protect old tasks but often limit new learning. (ii) Memory-replay strategies (Deep-Generative-Replay; iCaRL; chaudhry2019continual) interleaves past and current data, yet pretraining data of modern open-source MLLMs is usually unavailable for post-training. (iii) Architecture-based techniques (Progressive-Neural-Networks; Hard-Attention-to-the-Task) assign task-specific modules, but their parameter overhead makes them impractical for large MLLMs. Fortunately, with the rise of RFT algorithms such as GRPO (shao2024deepseekmath), recent studies (Seg-Zero; Jigsaw-R1; lai2025reinforcement) have shown that RFT can significantly reduce CF in MLLMs, although their analysis and explanations are still limited. Recently, RL’s Razor (Razor) argues that online RL mitigates CF because it is implicitly biased toward KL-minimal solutions. Aligning with this discovery, we further analyze why online sampling distribution inherently reduces forgetting from a data-centric perspective with learning dynamics theory. Besides, we would like to clarify that our paper isn’t aim to design a better algorithm than classical methods for continuous learning.
@@ 3.2 Rule-based Rewards and RFT
RFT Algorithm. We adopt Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) as our RFT algorithm. Formally, we maximize the following objective:
@@ 4 Experimental Setup
Dataset Construction. Jigsaw puzzles are built upon COCO 2014 (COCO) image dataset. For training, we sample around 22k images from the COCO training set to generate jigsaw puzzles. For testing, we sample images from the COCO test set. For the SFT dataset, we provide two data formats, i.e., Non-Reasoning data and Reasoning data: the first directly provides the ground-truth answers without reasoning processes, while the second additionally consists of reasoning trajectories generated by GPT-4o with both the question and the answer as input, dubbed as Rea-4o-Rollout.
MLLMs. We employ Qwen2.5-VL-3B (Qwen2.5-VL) and Qwen2.5-VL-7B as our MLLMs due to their strong performance on vision-language understanding and support of native resolution input.
@@
Grounding. RefCOCO/+/g (RefCOCO; RefCOCOg) test referring-expression comprehension, requiring the model to localize objects described by the free-form text.
OCR, chart & document understanding. DocVQA (DocVQA), InfoVQA (InfoVQA), and OCRBench (OCRBench) probe the ability of MLLMs to read and reason over scanned documents, forms, and scientific plots.
General VQA. MME (MME), MMStar (MMStar), and GQA (GQA) cover visual reasoning, spatial relations, and multimodal commonsense.
Hallucination. POPE (POPE) measures tendency of generation not grounded in image.
College-level Problems. MMMU (Mmmu) is a college-level multimodal benchmark spanning six disciplines for knowledge-grounded visual reasoning.
@@ 5.1 Can RFT master the novel jigsaw puzzles?
| Qwen2.5-VL-3B | Qwen2.5-VL-7B | |||||||||||||||||
| Base | RFT | SFT-Non-Rea |
Base | RFT | SFT-Non-Rea |
| Jigsaw Puzzles (test) | ||||||||||||||||||
| Training steps | – | 27,360 | 200 | 4,100 | 2,670 | – | 27,360 | 400 | 4,100 | 3,000 | ||||||||
| 33 puzzles | 0.0 | 66 (66) | 53.0 (53) | 70.0 (70) | 70.0 (70) | 0.0 | 75 (75) | 80 (80) | 78 (78) | 81 (81) | ||||||||
| Grounding | ||||||||||||||||||
| RefCOCO | 88.8 | 88.4 (0.4) | 6.1 (82.8) | 74.2 (14.6) | 84.6 (4.2) | 90.0 | 89.4 (0.6) | 32.9 (57.2) | 52.5 (37.5) | 81.4 (8.6) | ||||||||
| RefCOCO+ | 82.0 | 82.2 (0.2) | 4.2 (77.7) | 68.3 (13.6) | 77.6 (4.4) | 84.7 | 83.6 (1.1) | 28.8 (55.9) | 47.8 (36.9) | 75.1 (9.6) | ||||||||
| RefCOCOg | 86.0 | 84.1 (1.9) | 5.6 (80.4) | 71.3 (14.7) | 80.8 (5.2) | 86.4 | 86.3 (0.1) | 30.1 (56.4) | 48.3 (38.1) | 76.1 (10.3) | ||||||||
| Document & OCR | ||||||||||||||||||
| DocVQA | 92.8 | 91.5 (1.3) | 81.6 (11.3) | 90.3 (2.5) | 89.8 (3.1) | 94.4 | 94.4 (0.0) | 67.1 (27.4) | 92.1 (2.3) | 93.5 (0.9) | ||||||||
| InfoVQA | 74.3 | 73.1 (1.2) | 62.6 (11.7) | 71.4 (2.8) | 70.7 (3.6) | 80.1 | 79.1 (1.0) | 44.6 (35.5) | 75.6 (4.5) | 77.3 (2.8) | ||||||||
| OCRBench | 79.3 | 77.1 (2.1) | 65.9 (13.4) | 69.4 (9.9) | 74.9 (4.4) | 83.4 | 83.4 (0.0) | 51.7 (31.7) | 80.5 (2.9) | 81.4 (2.0) | ||||||||
| General VQA | ||||||||||||||||||
| MME | 2140 | 2137 (3) | 1631 (509) | 1478 (662) | 2132 (8) | 2333 | 2325 (8) | 479(1854) | 2084 (249) | 2207 (126) | ||||||||
| MMStar | 56.2 | 55.8 (0.5) | 49.2 (7.0) | 51.7 (4.5) | 52.2 (4.0) | 62.8 | 64.4 (1.7) | 0.0 (62.8) | 59.1 (3.7) | 60.4 (2.4) | ||||||||
| GQA | 60.1 | 59.5 (0.6) | 54.7 (5.4) | 50.0 (10.1) | 54.0 (6.1) | 60.4 | 60.3 (0.1) | 21.7 (38.7) | 53.5 (6.9) | 57.0 (3.3) | ||||||||
| Hallucination | ||||||||||||||||||
| POPE | 86.9 | 86.5 (0.3) | 85.9 (1.0) | 69.4 (17.5) | 85.4 (1.4) | 86.2 | 86.0 (0.2) | 16.3 (69.9) | 74.1 (12.1) | 83.1 (3.1) | ||||||||
| College-level Problems | ||||||||||||||||||
| MMMU | 46.9 | 46.3 (0.6) | 43.4 (3.5) | 43.0 (3.9) | 44.3 (2.6) | 51.3 | 50.0 (1.3) | 22.4 (28.9) | 46.1 (5.2) | 48.8 (2.6) | ||||||||
@@
We then examine whether reinforcement fine-tuning can enable the base model to learn entirely new tasks or knowledge from scratch. Specifically, we apply GRPO to Qwen-2.5-VL-3B/7B on the training set of jigsaw puzzles for 10 epochs, encouraging a comprehensive and sufficient exploration of the novel task. After convergence, the final models achieve an accuracy of 66%/75% on the held-out test set as shown in Tab. 1, dramatically outperforming the base model. Qualitative results on test examples, as in Fig. 11 of the Appendix, show that the model learns to generate meaningful reasoning processes before giving the final answers. Although prior work (yue2025does) suggests that RFT fails to induce fundamentally new reasoning patterns in base models, we show that with sufficiently long-term exploration, RFT can in fact enable the model to solve novel jigsaw puzzles from scratch.

### v4

# Why Reinforcement Fine-Tuning Preserves Prior Knowledge Better: A Data Perspective

@@ Abstract
Post-training algorithms such as Supervised Fine-Tuning (SFT) and Reinforcement Fine-Tuning (RFT) are widely used to adapt (multimodal) large language models to downstream tasks. While effective at task adaptation, their impact on retaining prior knowledge remains unclear. In this paper, we introduce jigsaw puzzles as a novel task absent from existing pretraining corpora and systematically study the behavior of SFT and RFT on the open-source Qwen2.5-VL series. Our experiments reveal a sharp trade-off: SFT enables rapid task acquisition but leads to catastrophic forgetting, whereas RFT learns more slowly but better maintains prior knowledge. We study this phenomenon through learning dynamics by examining both the magnitude and direction of how training data influence prior knowledge. Our analysis shows that RFT mainly reinforces correct samples naturally aligned with the base model’s probability landscape, leading to weaker interference with prior knowledge. Moreover, training on RFT-simulated rollouts, which exert a smaller magnitude of influence and are better aligned in direction to prior knowledge, allows SFT to preserve prior knowledge better while rapidly learning new tasks. We further validate our framework on Qwen2.5 post-training in math and scientific QA, observing consistent forgetting and learning-dynamics trends. These findings suggest that the distribution of post-training data, rather than algorithmic differences alone, plays a central role in forgetting, and highlight RFT as a promising ingredient for stable continual post-training.
@@ 1 Introduction
In the era of large models, two primary post-training methods, i.e., Supervised Fine-Tuning (SFT) (Wei et al., 2021) and Reinforcement Fine-Tuning (RFT) (DeepSeek-AI et al., 2025; Ouyang et al., 2022), have emerged for enhancing model performance on domain-specific tasks. These methods have been pivotal in enabling (multimodal) large language models to learn specific downstream tasks, follow human instructions, and acquire reasoning capabilities, yielding impressive results. However, current practices primarily focus on performance improvement for specific downstream tasks, while overlooking potential impact of fine-tuning algorithms on model’s pre-existing knowledge. This oversight raises concerns about the model’s ability to retain and apply prior knowledge.
@@
To this end, this paper investigates how post-training algorithms, specifically SFT and RFT, affect the retention of prior knowledge when large models are trained to learn entirely novel knowledge or tasks that were absent during pretraining. To establish a challenging and genuinely novel task for testing, we introduce jigsaw puzzles as the target task for learning, as in Fig. 1. Through preliminary experiments, we observe that existing state-of-the-art MLLMs, including GPT-4o (Hurst et al., 2024), fail to solve even simple 3x3 jigsaw puzzles, indicating that this task represents a novel problem not covered by current pretraining corpora. Thus, jigsaw puzzles can serve as a fair and meaningful task for evaluating impact of post-training algorithms on prior knowledge.
@@
We conduct systematic fine-tuning experiments using both standard SFT and RFT, i.e., GRPO (Shao et al., 2024), on open-sourced Qwen2.5-VL (Bai et al., 2025) series. Interestingly, we find that SFT can master novel tasks with solely hundreds or thousands of training steps, while RFT requires several tens of thousands of training steps to successfully solve jigsaw puzzles and achieves similar accuracy as SFT. This finding suggests that large-scale RFT can teach model to solve tasks that base model is completely unable to handle. In a sense, this shows that RFT can push the model beyond its original capability boundary. In addition to performance on novel tasks, we observe that SFT incurs severe forgetting of previous knowledge, with substantial performance drops across diverse benchmarks, especially on tasks with similar output formats as jigsaw puzzles. In contrast, RFT, by leveraging reward-driven credit assignments for the simulated rollouts, can master the novel jigsaw puzzles while maintaining decent performance on prior tasks.
@@
The distinct phenomenon between SFT and RFT naturally raises a question: Why SFT incurs catastrophic forgetting while RFT does not? While both algorithms increase likelihood of correct responses, RFT adaptively reweights likelihood of rollout with reward, whereas SFT uniformly increases likelihood of static human annotations. Inspired by this, we collect correct rollouts during RFT and use them as supervised data for SFT. Surprisingly, SFT trained on these correct rollouts not only acquires novel knowledge quickly but also preserves prior knowledge better. This suggests that construction of fine-tuning data, rather than training algorithm, is key factor in knowledge forgetting.
@@
Furthermore, we provide a new perspective based on learning dynamics (Ren and Sutherland, 2024), which links the likelihood change of prior knowledge to the gradient induced by an individual training example , on understanding this distinct forgetting behavior by analyzing the magnitude and direction of how training data influence prior knowledge. We first observe that SFT data without reasoning trajectories usually interfere more with prior knowledge, as verified with a much larger norm of empirical neural tangent kernel (eNTK) between SFT data and prior knowledge. While datasets with reasoning trajectories, such as reasoning trajectories generated by GPT-4o and collected during RFT, usually exhibit a smaller norm of eNTK and less forgetting of prior knowledge, implying that introducing explicit reasoning can help alleviate knowledge forgetting.
@@
As for the direction of interference, we find that data with reasoning trajectories generated by GPT-4o typically belong to high-perplexity regions of the base model. In contrast, data collected during RFT are naturally generated from regions where the base model already assigns a moderate probability. This suggests that pretraining has already shaped certain linguistic regions by accident that are well-aligned with novel tasks, while remaining compatible with prior knowledge. Importantly, according to learning dynamics, the influence of training on one example over the likelihood of another example is symmetric: increasing the likelihood of has the same marginal effect on as increasing has on . Therefore, when RFT discovers and reinforces such hidden linguistic regions shaped during pre-training, it degrades less the likelihood of prior knowledge . Crucially, such regions are difficult to identify during the stage of dataset construction for SFT, but are accessible through RFT’s active exploration within linguistic space. This highlights RFT as an effective algorithm for stable novel knowledge acquisition without suffering from catastrophic forgetting. Further, more experiments on Qwen2.5 post-training in math and scientific QA show consistent forgetting and learning-dynamics results. Formally, our contributions are three-fold:
@@
We show that RFT can solve unseen tasks while preserving prior competencies. Moreover, SFT trained on RFT-generated rollouts can match RFT’s performance while markedly reducing catastrophic forgetting, underscoring the central role of data construction in post-training.
@@ 2 Related Works
Jigsaw Puzzles. Jigsaw puzzles has long been a popular self-supervised task in the computer vision community, aimed at learning visual representations (Noroozi and Favaro, 2016; Carlucci et al., 2019) by spatial reasoning and part-whole understanding. Recently, this task has been repurposed for probing weak spot of MLLMs: Lyu et al. (2025) shows that leading MLLMs perform far behind than human performance. The contemporary work Jigsaw-R1 (Wang et al., 2025) solves jigsaw puzzles with RFT, achieving much better performance. Collectively, these works mainly treat jigsaw puzzles as pretext task for representation learning or test benchmark for MLLMs. However, we employ jigsaw puzzles to investigate how post-training algorithms affect the forgetting behavior of MLLMs.
@@
Reinforcement Fine-Tuning in MLLMs. Inspired by the success of RFT in large language models (DeepSeek-AI et al., 2025; Ouyang et al., 2022), recent work has applied RFT to MLLMs. Among them, Meng et al. (2025) finds that RFT can achieve better out-of-distribution generalization performance than SFT. Meanwhile, RFT is also employed for perception-centric tasks (Liu et al., 2025c; Shen et al., 2025; Liu et al., 2025b), still conferring notable gains in generalization and robustness. Concurrently, Jigsaw-R1 (Wang et al., 2025) introduced RFT to the novel task of jigsaw puzzles but achieved limited accuracy. Building on this direction, we extend RFT training to tens of thousands of steps to enable deeper exploration, yielding substantial gains on jigsaw puzzles.
@@
Catastrophic Forgetting. Early work (McCloskey and Cohen, 1989; Ratcliff, 1990) showed that even minimal sequential training on disjoint data can cause rapid “catastrophic forgetting” (CF). Existing strategies to mitigate CF fall into three categories: (i) Regularization-based methods (Kirkpatrick et al., 2016; Zenke et al., 2017; Li and Hoiem, 2017) constrain updates to protect old tasks but often limit new learning. (ii) Memory-replay strategies (Shin et al., 2017; Rebuffi et al., 2017; Chaudhry et al., 2019) interleaves past and current data, yet pretraining data of modern open-source MLLMs is usually unavailable for post-training. (iii) Architecture-based techniques (Rusu et al., 2016; Serrà et al., 2018) assign task-specific modules, but their parameter overhead makes them impractical for large MLLMs. Fortunately, with the rise of RFT algorithms such as GRPO (Shao et al., 2024), recent studies (Liu et al., 2025b; Wang et al., 2025; Lai et al., 2025) have shown that RFT can significantly reduce CF in MLLMs, although their analysis and explanations are still limited. Recently, RL’s Razor (Shenfeld et al., 2025) argues that online RL mitigates CF because it is implicitly biased toward KL-minimal solutions. Aligning with this discovery, we further analyze why online sampling distribution inherently reduces forgetting from a data-centric perspective with learning dynamics theory. Besides, we would like to clarify that our paper isn’t aim to design a better algorithm than classical methods for continuous learning.
@@ 3.2 Rule-based Rewards and RFT
RFT Algorithm. We adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as our RFT algorithm. Formally, we maximize the following objective:
@@ 4 Experimental Setup
Dataset Construction. Jigsaw puzzles are built upon COCO 2014 (Lin et al., 2014) image dataset. For training, we sample around 22k images from the COCO training set to generate jigsaw puzzles. For testing, we sample images from the COCO test set. For the SFT dataset, we provide two data formats, i.e., Non-Reasoning data and Reasoning data: the first directly provides the ground-truth answers without reasoning processes, while the second additionally consists of reasoning trajectories generated by GPT-4o with both the question and the answer as input, dubbed as Rea-4o-Rollout.
MLLMs. We employ Qwen2.5-VL-3B (Bai et al., 2025) and Qwen2.5-VL-7B as our MLLMs due to their strong performance on vision-language understanding and support of native resolution input.
@@ 5.1 Can RFT master the novel jigsaw puzzles?
| Qwen2.5-VL-3B | Qwen2.5-VL-7B | |||||||||
| Base | RFT | SFT-Non-Rea | SFT-Rea- 4o-Rollout | SFT-Rea- GRPO-Rollout | Base | RFT | SFT-Non-Rea | SFT-Rea- 4o-Rollout | SFT-Rea- GRPO-Rollout | |
| Jigsaw Puzzles (test) | ||||||||||
| Training steps | – | 27,360 | 200 | 4,100 | 2,670 | – | 27,360 | 400 | 4,100 | 3,000 |
| 33 puzzles | 0.0 | 66 (66) | 53.0 (53) | 70.0 (70) | 70.0 (70) | 0.0 | 75 (75) | 80 (80) | 78 (78) | 81 (81) |
| Grounding | ||||||||||
| RefCOCO | 88.8 | 88.4 (0.4) | 6.1 (82.8) | 74.2 (14.6) | 84.6 (4.2) | 90.0 | 89.4 (0.6) | 32.9 (57.2) | 52.5 (37.5) | 81.4 (8.6) |
| RefCOCO+ | 82.0 | 82.2 (0.2) | 4.2 (77.7) | 68.3 (13.6) | 77.6 (4.4) | 84.7 | 83.6 (1.1) | 28.8 (55.9) | 47.8 (36.9) | 75.1 (9.6) |
| RefCOCOg | 86.0 | 84.1 (1.9) | 5.6 (80.4) | 71.3 (14.7) | 80.8 (5.2) | 86.4 | 86.3 (0.1) | 30.1 (56.4) | 48.3 (38.1) | 76.1 (10.3) |
| Document & OCR | ||||||||||
| DocVQA | 92.8 | 91.5 (1.3) | 81.6 (11.3) | 90.3 (2.5) | 89.8 (3.1) | 94.4 | 94.4 (0.0) | 67.1 (27.4) | 92.1 (2.3) | 93.5 (0.9) |
| InfoVQA | 74.3 | 73.1 (1.2) | 62.6 (11.7) | 71.4 (2.8) | 70.7 (3.6) | 80.1 | 79.1 (1.0) | 44.6 (35.5) | 75.6 (4.5) | 77.3 (2.8) |
| OCRBench | 79.3 | 77.1 (2.1) | 65.9 (13.4) | 69.4 (9.9) | 74.9 (4.4) | 83.4 | 83.4 (0.0) | 51.7 (31.7) | 80.5 (2.9) | 81.4 (2.0) |
| General VQA | ||||||||||
| MME | 2140 | 2137 (3) | 1631 (509) | 1478 (662) | 2132 (8) | 2333 | 2325 (8) | 479(1854) | 2084 (249) | 2207 (126) |
| MMStar | 56.2 | 55.8 (0.5) | 49.2 (7.0) | 51.7 (4.5) | 52.2 (4.0) | 62.8 | 64.4 (1.7) | 0.0 (62.8) | 59.1 (3.7) | 60.4 (2.4) |
| GQA | 60.1 | 59.5 (0.6) | 54.7 (5.4) | 50.0 (10.1) | 54.0 (6.1) | 60.4 | 60.3 (0.1) | 21.7 (38.7) | 53.5 (6.9) | 57.0 (3.3) |
| Hallucination | ||||||||||
| POPE | 86.9 | 86.5 (0.3) | 85.9 (1.0) | 69.4 (17.5) | 85.4 (1.4) | 86.2 | 86.0 (0.2) | 16.3 (69.9) | 74.1 (12.1) | 83.1 (3.1) |
| College-level Problems | ||||||||||
| MMMU | 46.9 | 46.3 (0.6) | 43.4 (3.5) | 43.0 (3.9) | 44.3 (2.6) | 51.3 | 50.0 (1.3) | 22.4 (28.9) | 46.1 (5.2) | 48.8 (2.6) |
@@
We then examine whether reinforcement fine-tuning can enable the base model to learn entirely new tasks or knowledge from scratch. Specifically, we apply GRPO to Qwen-2.5-VL-3B/7B on the training set of jigsaw puzzles for 10 epochs, encouraging a comprehensive and sufficient exploration of the novel task. After convergence, the final models achieve an accuracy of 66%/75% on the held-out test set as shown in Tab. 1, dramatically outperforming the base model. Qualitative results on test examples, as in Fig. 11 of the Appendix, show that the model learns to generate meaningful reasoning processes before giving the final answers. Although prior work (Yue et al., 2025) suggests that RFT fails to induce fundamentally new reasoning patterns in base models, we show that with sufficiently long-term exploration, RFT can in fact enable the model to solve novel jigsaw puzzles from scratch.